A boat takes 3.3 hours to travel 10 km down a river, then 5.9 hours to return. You may want to review ( pages 876 - 878) . Part A Determine the speed of the water in the river.
Answer #1

Subtitute "r" in eq in 1
Answer #2

SOLUTION :


Let speed of water in the river be w km/hr and speed of boat be b km/hr


So, 


time taken to travel 10 km down the river = 10 / (b + w)

=> 3.3 = 10 / (b + w)

=> b + w = 10 / 3.3 = 3.03 km/hr ………… (1)


time taken to travel back 10 km up the river = 10 / (b - w)

=> 5.9 = 10 / (b - w)

=> b - w = 10 / 5.9  = 1,69 km/hr ………… (2)


Add (1) and (2) :

=> 2 b = 4.72 km/hr

=> b = 2.36 km/hr

=> w = 3.03 - b = 3.03 - 2.36 = 0.67 km/hr (ANSWER).

answered by: Tulsiram Garg